OpenStudy (anonymous):

|x/3 + 1| = 6 Solve. There should be a solution set. |x/3 + 1| = 6 Solve. There should be a solution set. @Mathematics

OpenStudy (amistre64):

when does: x/3 + 1 = 6 and x/3 + 1 = -6 ??

OpenStudy (amistre64):

|6| = 6 |-6| = 6 so there are only 2 options to satisfy

OpenStudy (anonymous):

Oh wow I totally forgot what you could plug in possible answers. Thanks.

OpenStudy (amistre64):

:) good luck

OpenStudy (anonymous):

How would I go about solving this equation for y: 4x + 3y = -12. ?

OpenStudy (amistre64):

subtract 4x from both sides and divide it all by 3

OpenStudy (amistre64):

4x + 3y = -12 -4x -4x ---------------- 3y = -4x-12 \3 \3 ------------ y = (-4/3)x - 4

OpenStudy (anonymous):

Thanks. This is the option that best matches the answer you gave me.. y = -4 - 4/3 x